# Trapped fireshell (halo) of photons and pairs around black-hole horizon: source for ultra-high-energy particles

### v1

@@ Abstract
We study the Compton-rocket effect of multi-photon interacting with electrons in an opaque fireball (or fire spot) of photons and $e^{-}e^{+}$ pairs at temperature $T_{\gamma}\gg m_{e}$. We find the charged-particle acceleration and the avalanche runaway process, leading to a non-trivial probability of ultra-high-energy (UHE) electrons and protons, which subsequently produce very-high-energy (VHE) photons and neutrinos. We show such peculiar dynamics using the Gamma-Ray Burst central engine fireball, whose inner part inflows and forms a gravitationally trapped fireshell (halo) around a black hole. The halo is a metastable, cooling via UHE particle emissions and blackbody radiation. We calculate the UHE particle luminosity varying in time, and discuss the peculiar features of such produced UHE particles, which lead to VHE particles, in connection with possible numerical simulations, observations and experiments.
@@ 1 Introduction
The discovery of ultra-high-energy (UHE $\sim 10^{18-20}$ eV) charged particles (cosmic rays) has been for decades. The acceleration mechanism of UHE charged particles remains mysterious, although much theoretical and observational progress has been made to understand their nature and identify their sources (see review [1]). Very-high-energy (VHE $\sim 10^{12-15}$ eV) neutral particles (photons and neutrinos) have been recently observed by the LHAASO [2], MAGIC [3] and ICECUB [4] collaborations. Much theoretical attention has been given to understanding the origin of VHE neutral particles. In principle, the primary UHE charged particles are energetic enough to produce secondary VHE neutral particles. Thus, two highly energetic particle events may correlate and originate from compact and powerful astrophysical sources, such as gravitational accretion disks, massive collapses, binary coalescence and active galactic nuclei. In these arenas, strong radiation fields and charged particles non-linearly interact via the Compton rocket effect [5, 6, 7] on relativistic electrons by the anisotropy of multi-photons and electron scattering. This dynamics has been studied by particle-in-cell numerical simulations [8, 9, 10, 11].
@@
One of these powerful astrophysical sources can be the central engine for Gamma-ray bursts (GRBs), an opaque fireballs of large ultra-relativistic particle energy density and pressure [12], and see reviews [13, 14, 15, 16, 17, 18, 19, 20]. Within the general relativity (GR) framework, we studied [21, 22] the hydrodynamical outflow of fireballs by using also the Boltzmann rate equation for electrons and positrons (pairs) annihilation to photons, obtaining the relation between the photon heat and baryon kinetic energies for main bursts and afterglow radiations. Further scrutiny study [23] shows in Fig. 1 that the fireball’s inner part is trapped due to strong gravitational attraction, prevailing over thermal pressure repulsion. On the contrary, its outer part outflows, accounting for the GRB phenomenon. The separatix between the inner and outer parts is $\bar{R}\approx 4GM=2r_{+}$ near the horizon.
@@
We will show in Section 2 that through the gravitational and thermal forces’ equilibrium, the trapped fireball’s inner part (halo around a black hole) forms a metastable thermal fluid of photons and pairs at high energy density and temperature. Using such an opaque horizon halo as an example, we study the Compton-rocket effect on accelerating charged particles (Section 3) and the avalanche runaway process (Section 4) leading to UHE electrons and protons. We calculate the luminosity of UHE particles and the halo cooling in Section 5, and discuss the peculiar features of UHE and VHE particles produced in this dynamics, in connection with numerical simulations, observations, and experiments in the final Section 6.
@@
Figure 1: Figure 2 of Ref. [23] shows the world lines of fireball hydrodynamics expanding outflow (dot line) and trapped inflow (solid line) in the Schwarzschild geometry of a collapsing core (dashed line) of the mass $M$, and horizon radius $r_{+}=2M$. The separatrix radius between out and in flows is $\bar{R}=4M$, and the trapped inflow energy is about half of the expanding outflow energy. We can approximately obtain the separatrix by estimating the balance between thermal repulsive pressure and gravitational attractive force $4\pi r_{+}^{2}\rho_{\gamma}(r_{+})c\sim Gm_{\gamma}M/r_{+}^{2}$ on the photon-pair shell of mass energy $m_{\gamma}=\rho_{\gamma}(r_{+})4\pi r_{+}^{2}\bar{R}$. It is in distinct contrast with the processes of pressure-less baryon dust or a test particle free-falling in a gravitational collapse. The gravitational time dilation $\delta t=\left(1-{\frac{2Gm(r)}{r}}\right)^{-1/2}\delta\tau\rightarrow\infty$ for an observer at infinity is evident near the horizon $r\rightarrow r_{+}$, and $\delta\tau$ is a finite proper time (distance) for a local observer.
@@ 2 Trapped fireshell (halo) around black hole horizon
The highly energetic and dense fireshell comprises ultra-relativistic particles and antiparticles, for example, photons $\gamma$, electrons $e^{-}$ and positrons $e^{+}$ pairs, undergoing the back and forth process $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$, which establishes a local thermal equilibrium at high temperatures. We describe the fireball as a perfect fluid of energy density $\rho_{\gamma}(r)$ and pressure $P_{\gamma}(r)$ with the Equation of State $P_{\gamma}(r)=\rho_{\gamma}(r)/3$}. Baryonic matter is ionised, and its amount and binding energy are negligible in studying fireballs at high temperatures.
@@ 2.1 Tolman–Oppenheimer–Volkoff equation for halo energy density
We adopt the spherically symmetric Tolman–Oppenheimer–Volkoff (TOV) equation to study the gravitational equilibrium configuration of ultra-relativistic photons and electron-positron pairs trapped by the gravitational force near the horizon,
@@
in the Schwarzschild geometry of radial coordinate $r$ and time $t$ of an observer at infinity. The TOV equation is the equilibrium balance between the gravitational and radiation forces. Within the $r$-sphere, the total energy mass ${\textstyle m(r)=M+E_{\gamma}(r)}\approx M$ and the photon-pair fluid energy $E_{\gamma}(r)$ is much smaller than the black hole mass $M$.
Figure 2: Left: the trapped fireshell (halo) energy density $\rho_{\gamma}$ is expressed in units of $m_{e}^{4}$ and as functions of radius $r$, which is in the unit of the horizon radius $r_{+}=2GM=2.33\times 10^{6}$ cm for the black hole mass $M=7.75M_{\odot}$ (2.3). Right: the trapped fireshell (halo) temperature $T_{\gamma}$ is expressed in units of the electron mass $m_{e}\approx 0.5{\rm MeV}$. The maximal temperature $T^{+}_{\gamma}$ near the horizon rapidly decreases to the asymptotic value $\approx 2m_{e}$ near the boundary $2r_{+}=4GM$, where the electron-positron pair annihilation takes place. The Compton energy density $m_{e}^{4}\approx 4.54\times 10^{21}{\rm erg/cm}^{3}$ is of the same order as the critical energy density $\rho_{c}=E_{c}^{2}/(8\pi)=5.45m_{e}^{4}$, where $E_{c}=m_{e}^{2}/e$ is the critical electric field for electron and positron pair production in the semi-classical approximation [19]. We use the $\hbar=c=1$ and Compton unit of the electron mass $m_{e}$, unless otherwise specified.
Rescaling the radius $r$ in the unit of $2GM$ and density $\rho_{\gamma}$ in the unit of $\rho_{{}_{M}}=M/V_{M}$, where “volume” $V_{M}=(4\pi/3)(2GM)^{3}$, we recast the TOV equation (2.1) as a dimensionless equation
@@
for the density $\rho_{\gamma}$. The photons and pairs flux $F_{\gamma}=\rho_{\gamma}c$ and luminosity $L_{\gamma}=4\pi r^{2}\rho_{\gamma}c$.
The “density” $\rho_{{}_{M}}=M/V_{M}$ varies from $(2-10)\rho_{n}$ for the mass $M$ varies from $(3-10)M_{\odot}$. Here, we consider a massive stellar baryon core of neutron stars at the nuclear density $\rho_{n}$, gravitationally collapsing to a compact object of size about $2GM$ of the black hole horizon. The average baryon density increases and becomes about $5\sim 8$ times the nuclear energy density. We choose (i) $\rho_{{}_{M}}=10\rho_{n}$ that effectively corresponds to a black hole of the mass $M$ and horizon radius $r_{+}$
$M=7.75M_{\odot},\quad r_{+}=2GM=2.33\times 10^{6}{\rm cm};$ (2.3)
(ii) the photon and pair density and luminosity near the horizon are
$\rho_{\gamma}^{+}\equiv\rho_{\gamma}(r_{+})=\eta\rho_{{}_{M}}=\rho_{n},~~~L_{\gamma}^{+}\equiv L_{\gamma}(r_{+})=\eta L_{\rm pl},$ (2.4)
$\eta=0.1$ and the “Planck luminosity” $L_{\rm pl}\equiv 4\pi r_{+}^{2}\rho_{{}_{M}}c=(3/2)m^{2}_{\rm pl}=3.73\times 10^{59}{\rm ergs/sec}$. These are the boundary conditions for the ordinary differential equation (2.2). We chose $\rho_{\gamma}^{+}=\rho_{n}$ for the reason that at high temperatures, many species of relativistic particles, lepton-antilepton pairs and photons, light quark-antiquark pairs and gluons, can be excited, contributing to the energy density $\rho_{\gamma}^{+}$. It is a fluid of lepton-photon and quark-gluon plasma. The essential quantity is the averaged energy over species $d_{f}$, i.e., the fluid temperature $T_{\gamma}(r)$ (2.8) and value $T^{+}_{\gamma}$ (2.9) near the horizon.
Solving the TOV equation (2.2) with boundary conditions (2.4), we show on the left (right) of Figs. 2 that the energy density $\rho_{\gamma}$ (the temperature $T_{\gamma}$) of photons and pairs is maximal near the horizon $r_{+}$ and rapidly decreases as the radius increases to $2r_{+}$. The “trapped fireshell (halo)” is named for such an energetic fluid configuration of photons and pairs near the black hole horizon, where the maximal gravity effect confines the opaque fluid of photons and pairs. The macroscopic time scale of hydrodynamical relaxation for establishing a trapped fireshell is about $\bar{R}/v_{s}$, where the fluid sound velocity $v_{s}=(P_{\gamma}/\rho_{\gamma})^{1/2}=c/\sqrt{3}$.
@@
$E_{\gamma}(r)$ $=$ $\int_{r_{+}}^{2r_{+}}\rho_{\gamma}(r)\frac{4\pi r^{2}dr}{(1-\frac{r_{+}}{r})^{1/2}}=\rho_{\gamma}^{+}4\pi r_{+}^{2}\delta\ell,$ (2.5)
which is the energy of photons and pairs, as will be discussed below. Since the energy density $\rho_{\gamma}(r)$ of photons and pairs is highly peaked near the horizon, as shown in Figs. 2, we use the $\delta$-function approximation $\int(1-r_{+}/r)^{-1/2}\rho_{\gamma}(r)dr\approx\rho_{\gamma}(r_{+})\delta\ell$, and $\delta\ell=(1-r_{+}/r)^{-1/2}dr$ is a proper distance near the horizon. The macroscopic length $\delta\ell$ should be much larger than the Compton length $\lambda_{e}$, but smaller than the halo size $2GM\approx 2.33\times 10^{6}$cm. The total halo energy $E_{\gamma}$ must be less than the maximal gravitational binding energy $GM^{2}/(2r)|_{r=2GM}=M/4$, which can be gained in collapses.
2.2 Thermodynamic and opaque fluid of photons and pairs
In the local Lorentz frame, we discuss thermodynamic quantities, e.g., temperature $T_{\gamma}$, of the opaque fluid of photons and pairs. The photons and electron-positron pairs’ energy (number) density $\rho_{\gamma}$ ($n_{\gamma}$) is extremely large inside the trapped fireshell. The mean free path $\xi_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$ and the thermalization time scale $\tau_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}c$ are much smaller than the macroscopic size $\bar{R}$ and the time scale $\bar{R}/v_{s}$. Therefore, the local thermal equilibrium is almost instantly established via the back-and-forth process $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$ of the cross section,

### v2

@@ Abstract
We study the Compton-rocket effect of strong radiation force accelerating electrons in an opaque fireshell (or fire spot) of dense photons and electron-positron pairs, whose temperature is spatially inhomogeneous and exceeds the electron mass. We find the possibility of the charged-particle acceleration and the avalanche runaway process, leading to a non-trivial probability of ultra-high-energy (UHE) electrons and protons, which subsequently produce very-high-energy (VHE) photons and neutrinos. In a simplified one-dimensional model, we qualitatively show such peculiar dynamics using the fireball, Gamma-Ray Burst central engine, whose inner part inflows and forms a gravitationally trapped fireshell (halo) around the horizon of a black hole. The fireshell is metastable, cooling via UHE particle emissions and blackbody radiation. We calculate the UHE particle luminosity varying in time, and discuss the peculiar features of such produced UHE particles, which lead to VHE particles, in connection with possible numerical simulations, observations and experiments.
@@ 1 Introduction
The discovery of ultra-high-energy (UHE $\sim 10^{18-20}$ eV) charged particles (cosmic rays) has been for decades. The acceleration mechanism of UHE charged particles remains mysterious, although much theoretical and observational progress has been made to understand their nature and identify their sources (see review [1]). Very-high-energy (VHE $\sim 10^{12-15}$ eV) neutral particles (photons and neutrinos) have been recently observed by the LHAASO [2], MAGIC [3], and IceCube [4] collaborations. Much theoretical attention has been given to understanding the physical origin of UHE charged particles and VHE neutral particles, and identifying astrophysical sources where they come from.
It is physically reasonable to assume that these highly energetic particles may originate from compact and powerful astrophysical sources, such as accretion disks, massive core collapses, binary coalescence and active galactic nuclei. In these sources, there may exist highly energetic arenas, where strong gravitational fields, dense radiation fields, and charged particles interact non-linearly with one another. These interactions possibly result in the high-energy processes of accelerating charged particles, leading to UHE and VHE particles.
@@
One of these powerful astrophysical sources can be the central engine for Gamma-ray bursts (GRBs), an opaque fireballs of large ultra-relativistic particle (photon and pair) energy density and pressure [5], and see reviews [6, 7, 8, 9, 10, 11, 12, 13]. Within the general relativity (GR) framework, we studied [14, 15] the hydrodynamical outflow of fireballs by using also the Boltzmann rate equation for electrons and positrons (pairs) annihilation to photons, obtaining the relation between the photon heat and baryon kinetic energies for main bursts and afterglow radiations. We will first use the fireball, Gamma-Ray Bursts (GRBs) central engine, to show a possible highly energetic arena, a fireshell or fire spot, then illustrate our studies and results of the high-energy interactions and processes that produce UHE and VHE particles in such an arena.
@@
We will study in Section 2 that through the gravitational force and thermal pressure, the fireball’s inner part (halo around a black hole) is trapped and forms a metastable thermal fluid of photons and pairs at high energy density and temperature. Using such an opaque horizon halo, we study the Compton-rocket effect on accelerating charged particles (Section 3) and the avalanche runaway process (Section 4) leading to UHE electrons and protons. We calculate the luminosity of UHE particles and the halo cooling in Section 6, and discuss the peculiar features of UHE and VHE particles produced in this dynamics, in connection with numerical simulations, observations, and experiments in the final Section 7. We use the $\hbar=c=1$ and Compton unit of the electron mass $m_{e}$, unless otherwise specified in necessary cases for clarification.
@@
Figure 1: Figure 2 of Ref. [16] shows the world lines of fireball hydrodynamics expanding outflow (dot line) and trapped inflow (solid line) in the Schwarzschild geometry of a collapsing massive stellar core (dashed line), which forms a black hole of mass $M$. Near the horizon radius $r_{+}=2M$ of a formed black hole, the gravitational time dilation $\delta t\gg\delta\tau$ for an observer at a distance is evident; $\delta\tau$ is a finite proper time (distance) for a local observer. The separatrix radius between out and in flows is $\bar{R}=4M$, and the trapped inflow energy is comparable with the expanding outflow energy. The total inner and outer fireball energy gain from the collapsing system’s gravitational binding energy, and it should be much smaller than $(M^{2}/r_{\infty}-M^{2}/r_{+})/2=M/4$ in the Newtonian approximation.
@@ 2 Trapped fireshell (halo) around black hole horizon
2.1 Trapped fireshell (halo) of energetic and dense photons and pairs
Although the dynamics of fireball formation are still under investigation  [^1] In Refs. [17, 18], we present the detailed studies of the possibility of how fireballs or jets form during gravitational collapses of massive rotating stellar cores or binary coalescence: heavy hadrons gain kinetic energy from gravitational potential, and their heavy-ion collisions produce photons and pairs, resulting in the fireball or jet formation., the hydrodynamical outflow of fireballs accounting for GRBs has been well understood upon the assumptions that entire fireballs are propelled out in a flat spacetime. Nevertheless, a scrutiny GR study [16] investigates the hydrodynamical evolution of fireballs in the Schwarzschild geometry of a massive stellar core, which undergoes a gravitational collapse process to form a black hole of mass $M$. The result shows in Fig. 1 that the fireball’s inner part $(r<\bar{R})$ is trapped due to strong gravitational attraction, prevailing over thermal pressure repulsion. On the contrary, its outer part $(r>\bar{R})$, where gravitational attraction is weak and thermal pressure repulsion prevails, hydrodynamically outflows, undergoing internal and external shocks in its evolution, accounting for the GRB phenomenon as usually described. The radius separatix between the inner and outer parts is $\bar{R}\approx 4GM=2r_{+}$ near the black hole horizon $r_{+}=2M$  [^2] We can approximately estimate the separatrix $\bar{R}$ based on the balance between thermal repulsive pressure and gravitational attractive force $4\pi r_{+}^{2}\rho_{\gamma}(r_{+})c\sim Gm_{\gamma}M/r_{+}^{2}$ on the fireball inner part of energy mass $m_{\gamma}\sim\rho_{\gamma}(r_{+})4\pi r_{+}^{2}\bar{R}$ and fireball density $\rho_{\gamma}(r_{+})$ at the horizon..
The gravitational collapsing system comprises a cold baryon core and a hot fireball of relativistic particles. When the spacetime horizon is approaching, the final phase of the gravitational collapse of the nonrelativistic baryon matter core and the relativistic particles’ fluid (fireball inner part), which becomes denser, must be nontrivial. It can neither be a free-fall process of the time scale $\sim GM/c$, nor a rapid transit process, like a GRB prompt emission of an expanding fireball at its transparency. In the final phase of gravitational collapses, the gravitational time dilation effect is significant for an observer at a distance, see Fig. 1 and its caption.
The trapped fireball’s inner part, which we call the fireshell (or halo) for short, is highly energetic and dense, comprising numerous ultra-relativistic particles and antiparticles, which are represented by photons $\gamma$, electrons $e^{-}$ and positrons $e^{+}$ pairs. They have zero chemical potential and are electrically neutral in total. It is completely opaque, undergoing the rapid back-and-forth process
$\gamma+\gamma\leftrightarrow e^{+}+e^{-},$ (2.1)
which establishes a local thermal equilibrium at high temperatures and densities. Since the spacetime length scales of the back-and-forth process (2.1) are very small, we describe all photons $\gamma$ and pairs $e^{+}e^{-}$ together as a radiation fluid, rather than identifying them as individual photons, electrons $e^{-}$ and positrons $e^{+}$  [^3] In other words, relativistic fermions always pair with their antiparticles via exchanging two photons. Therefore, we ignore the notion of anti-particles like positrons $e^{+}$ in this article.. We treat a local element of the trapped fireball as a perfect radiation fluid of large energy density $\rho_{\gamma}(r)$ and pressure $P_{\gamma}(r)$ with the Equation of State $P_{\gamma}(r)=\rho_{\gamma}(r)/3$. They are inhomogeneous as functions of radius $r$, due to the presence of the gravitational field of a massive collapsing stellar core or black hole. Once more, we stress that the total thermal energy of the trapped fireball’s inner part is comparable to that of the expanding fireball’s outer part for GRBs [16].
Moreover, in distinct contrast with the gravitational collapse processes of pressure-less baryon dust, the dynamics of inflow contraction of the opaque fireshell is much more complex, because the gravitational potential, the fireshell density and pressure increase. To study such an inflow contraction in time, we should, in principle, consider the inhomogeneous radiative fluid and Einstein equation in the spherically symmetric Lemaître spacetime [19]  [^4] The Lemaître metric reduces to the Lemaître-Tolman-Bondi one in the pressureless dust case, whose solutions are well known. However, despite many years of work, the solution of the Lemaître metric with a general perfect fluid source is not known. in the presence of a black hole, whose metric elements are functions of the coordinate radius $r$ and time $t$. Observe that (i) the fireshell inflow contraction should be slow in time because of its pressure being very large; (ii) the fireshell is deeply opaque so that its thermal energy dissipation rate is small. Therefore, the fireshell in Lemaître spacetime should be a metastable configuration undergoing slow contraction adiabatically. Namely, the Lemaître metric elements, the fireshell energy density and pressure should be slowly varying in time. Thus, we approximately adopt the static limit, all quantities being independent of time, the Lemaître spacetime reduces to the Tolman–Oppenheimer–Volkoff (TOV) one [20] in the interior of the fireshell with the exterior in Schwarzschild coordinates.
@@ 2.2 Tolman–Oppenheimer–Volkoff equation for halo energy density
In this section, we adopt the spherically symmetric TOV equation for the fireshell to study the gravitational equilibrium configuration of ultra-relativistic photons and pairs trapped in the Schwarzschild spacetime near the horizon,
@@
in the Schwarzschild geometry of radial coordinate $r$ and time $t$ of an observer at infinity. The TOV equation describes the equilibrium balance between the gravitational and radiation forces. Within the sphere of a radius $r$, the total energy mass ${\textstyle m(r)=M+E_{\gamma}(r)}\approx M$ and the photon-pair fluid energy $E_{\gamma}(r)$ is much smaller than the black hole mass $M$.
Figure 2: Left: the trapped fireshell (halo) energy density $\rho_{\gamma}$, solution to (2.2), is expressed in units of $m_{e}^{4}$ and as functions of radius $r$, which is in the unit of the horizon radius $r_{+}=2GM=2.33\times 10^{6}$ cm for the black hole mass $M=7.75M_{\odot}$ (2.4). Right: the trapped fireshell (halo) temperature $T_{\gamma}=(\rho_{\gamma}/\sigma_{{}_{\rm SB}}d_{f})^{1/4}$ (2.9) is expressed in units of the electron mass $m_{e}\approx 0.5{\rm MeV}$. The maximal temperature $T^{+}_{\gamma}$ near the horizon rapidly decreases to the asymptotic value $\approx 2m_{e}$ near the boundary $2r_{+}=4GM$, where the electron-positron pair annihilation takes place. The Compton energy density $m_{e}^{4}\approx 4.54\times 10^{21}{\rm erg/cm}^{3}$ is of the same order as the critical energy density $\rho_{c}=E_{c}^{2}/(8\pi)=5.45m_{e}^{4}$, where $E_{c}=m_{e}^{2}/e$ is the critical electric field for electron and positron pair production in the semi-classical approximation [12].
Rescaling the radius $r$ in the unit of $2GM$ and density $\rho_{\gamma}$ in the unit of “average density” $\rho_{{}_{M}}=M/V_{M}$, where “volume” $V_{M}=(4\pi/3)(2GM)^{3}$, we recast the TOV equation (2.2) as a dimensionless equation
@@
for the density $\rho_{\gamma}$. We consider a massive stellar baryon core of neutron stars at the nuclear saturation density $\rho_{n}\approx 0.16\,{\rm fm}^{-3}=1.6\times 10^{38}/{\rm cm}^{3}$, gravitationally collapsing to a compact object of size about $2GM$ of the black hole horizon. The “average density” $\rho_{{}_{M}}=M/V_{M}$ varies from $(2-10)\rho_{n}$ for the mass $M$ varying from $(3-10)M_{\odot}$. The value $\rho_{{}_{M}}=10\rho_{n}$ effectively corresponds to a black hole of the mass $M$ and horizon radius $r_{+}$
$M=7.75M_{\odot},\quad r_{+}=2GM=2.33\times 10^{6}{\rm cm}.$ (2.4)
The photon-pair pressure $P_{\gamma}(r)=\rho_{\gamma}(r)/3$ should be maximum near the horizon, where the gravitational force is the strongest. To illustrate the fireshell properties, we choose the photon-pair density near the horizon
$\rho_{\gamma}^{+}\equiv\rho_{\gamma}(r_{+})=\eta\rho_{{}_{M}}=\rho_{n},~~~\eta=0.1,$ (2.5)
as the boundary conditions for the TOV equation (2.3). The reason for $\rho_{\gamma}^{+}\sim\rho_{n}$ is that at high energies, many species of relativistic particles, lepton-antilepton pairs and photons, light quark-antiquark pairs and gluons, could be excited, contributing to the energy density $\rho_{\gamma}^{+}$. It turns out to be a fluid of lepton(pair)-photon and quark(pair)-gluon plasma. For the sake of simplified notations, we use the photon ($\gamma$) and pair ($e^{-}e^{+}$) notations representing all possible gauge bosons and fermion-antifermion pairs, and use the subscript $\gamma$” for the quantities of the trapped fireshell of photons and pairs.
Solving the TOV equation (2.3) with boundary conditions (2.5), we show in Fig. 2 (left) that the energy density $\rho_{\gamma}$ of photons and pairs is maximal near the horizon $r_{+}$ and rapidly decreases as the radius increases to $2r_{+}$. The “trapped fireshell (halo)” is named for such an energetic radiation fluid configuration of photons and pairs near the black hole horizon, where the maximal gravity effect confines the opaque radiation fluid of photons and pairs at high densities. The macroscopic time scale of hydrodynamical relaxation for establishing a trapped fireshell is about $\bar{R}/v_{s}$, where the fluid sound velocity $v_{s}=(P_{\gamma}/\rho_{\gamma})^{1/2}=c/\sqrt{3}$.
@@
$E_{\gamma}$ $=$ $\int_{r_{+}}^{2r_{+}}\rho_{\gamma}(r)\frac{4\pi r^{2}dr}{(1-\frac{r_{+}}{r})^{1/2}}=\rho_{\gamma}^{+}4\pi r_{+}^{2}\delta\ell,$ (2.6)
which is the thermal energy of photons and pairs, as will be discussed in the next section. Note that the trapped thermal energy is not small, but comparable with the expanding outflow energy for GRBs, see Fig. 2 caption and Ref. [16]. Since the energy density $\rho_{\gamma}(r)$ of photons and pairs is highly peaked near the horizon, as shown in Fig. 2, we use the $\delta$-function approximation $\int(1-r_{+}/r)^{-1/2}\rho_{\gamma}(r)dr\approx\rho_{\gamma}(r_{+})\delta\ell$, and $\delta\ell=(1-r_{+}/r)^{-1/2}dr$ is a proper distance near the horizon. The macroscopic length $\delta\ell$ should be much larger than the Compton length $\lambda_{e}$, but smaller than the fireshell size $2GM\approx 2.33\times 10^{6}$cm.
2.3 Thermodynamic and opaque radiation fluid of photons and pairs
The photons and pairs’ energy (number) density $\rho_{\gamma}$ ($n_{\gamma}$) is extremely large inside the trapped fireshell. The mean free path $\xi_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$ and the thermalization time scale $\tau_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}c$ are much smaller than the macroscopic fireshell size $\bar{R}$ and the time scale $\bar{R}/v_{s}$. Therefore, the local thermal equilibrium is almost instantly established via the back-and-forth process $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$ (2.1) of the cross section,